# KAN-ODEs: Kolmogorov-Arnold Network Ordinary Differential Equations for Learning Dynamical Systems and Hidden Physics

### v1

@@ III Experiments
We present three applications in this work to demonstrate the capability of KAN-ODEs. We validate our methodology and compare its performance against the standard MLP-based Neural ODE in Sec. III.1, where the dynamics of the Lotka-Volterra predator-prey model are inferred. Then, we tackle a one-dimensional PDE in Sec. III.2, using a KAN-ODE to infer a symbolic source term in a wave propagation simulation. This example demonstrates the flexibility of KAN-ODEs to be used as submodels in higher complexity simulations, as well as their capability to extract symbolic models from data. Finally, we learn the dynamics of the one-dimensional Burgers’ equation via a KAN-ODE to demonstrate its larger-scale capability as a standalone solver.
III.1 KAN-ODEs vs Neural ODEs: Lotka-Volterra
In this section, we target the Lotka-Volterra predator-prey model shown in Eq. 11. We replicate the simulation parameters used in the Neural ODE study of [3]. Namely, we use $\alpha=1.5$, $\beta=1$, $\gamma=1$, and $\delta=3$ with an initial condition of $\mathbf{u}_{0}=[x_{0},y_{0}]=[1,1]$ and a time span of $t\in[0,14]$ s. The first 3.5 s are used to train the KAN-ODE, while the remainder of the time history is withheld to validate the KAN-ODE’s ability to extrapolate to unseen times. The temporal grid used for training and loss evaluation has a spacing of 0.1s. The inputs to each KAN layer in this example are normalized to be on the $[-1,1]$ range required by the RBF networks via the hyperbolic tangent function, as in [22]. Training data is shown in Fig. 3(A).
@@ III.1 KAN-ODEs vs Neural ODEs: Lotka-Volterra
Loss profiles from training a 240-parameter KAN-ODE of shape [2, 10, 5], [10, 2, 5] are shown in Fig. 3(B1). Strong convergence is observed as early as $2\times 10^{4}$ epochs into training, while convergence into the $10^{-7}$ range of MSE loss occurs toward $10^{5}$ epochs. On a single CPU core, training a KAN of this size takes roughly 20 minutes per $10^{4}$ epochs. The converged KAN-ODE prediction is shown in Fig. 3(A), where the agreement in the testing window (t=3.5:14) is nearly indistinguishable from the agreement in the training window (t=0:3.5), and both exhibit excellent adherence to the synthetic data. We remark again that the trained KAN-ODE is reconstructing both the seen training data and unseen testing data from only the initial condition ($\mathbf{u}_{0}=[1,1]$), and can deliver output values on any grid and at any time, thanks to its coupling to an ODE solver.
Figure 3: Comparison between KAN-ODE and Neural ODE for L-V predator-prey model. (A) Synthetic data (training and testing) and KAN-ODE reconstruction. (B) Loss profile during training for (B1) KAN-ODE and (B2) MLP-based Neural ODE of comparable sizes (240 and 252, respectively). (C) Comparison of converged KAN-ODE and Neural ODE error using different model sizes, and two MLP depths (d=2 and d=3). Neural scaling rates of $N^{-2}$ and $N^{-4}$ plotted for comparison, as per the theory in [21]. (D) KAN of 4 hidden layers used to learn L-V dynamics (80 parameters)
Table 1: Lotka-Volterra Network Sizes and Performance
@@
To benchmark the KAN-ODE performance, testing was also carried out using a standard, MLP-based Neural ODE. In this case, the network size and training procedure were taken from the Neural ODE work of [3], to avoid biasing the results in the KAN-ODE’s favor. Namely, an MLP with a hidden layer comprising 50 nodes and the hyperbolic tangent activation function replaced the KAN in Eq. 7. Other relevant training details such as the train/test time windows, temporal grid, Lotka-Volterra equation parameters, and ODE solver remain identical to the KAN-ODE example, which itself was modeled as directly as possible after the Neural ODE efforts of [3]. The MLP defined here contains 252 trainable parameters (refer to the bold architecture in Table 1), which is comparable (and in fact slightly larger) than the 240 used in the reported KAN-ODE.
The Neural ODE iterates roughly three times faster than the similarly-sized KAN-ODE, taking closer to 7 minutes per $10^{4}$ epochs on the same single CPU core. However, it achieves a minimum of $3\times 10^{-5}$ training error over $10^{5}$ epochs, which is not as low as the $2.6\times 10^{-5}$ achieved by the KAN-ODE in just $10^{4}$ epochs, and nearly two orders of magnitude worse than the $8.3\times 10^{-7}$ training error achieved by the KAN-ODE in the same $10^{5}$ epochs. We remark that $3\times 10^{-5}$ is already a remarkably low amount of training error, and in fact do not add Neural ODE profiles to the plot in Fig. 3(A) as they would be nearly impossible to distinguish from the KAN-ODE profiles. We retain discussion of loss metrics, however, as the distinctions and comparisons made here are likely to be indicative of similar performance metrics in larger-scale, more complex problems that are not as straightforward to fit with such strong accuracy.
@@
The Neural ODE model qualitatively appears to overfit less than the KAN-ODE, especially in the later epochs. Quantitatively, however, the KAN-ODE still beats the MLP-based Neural ODE in accuracy to the unseen testing data. Even in the last 10% of the training cycle, i.e. epochs $90,000-100,000$, where the KAN-ODE appears to overfit substantially, it still achieves an average testing loss of $6.8\times 10^{-3}$, with a minimum value in that range of $1.9\times 10^{-5}$. The Neural ODE in this same range averages a much larger $1.4\times 10^{-2}$, with a minimum value of $5.4\times 10^{-5}$. If training for the KAN-ODE is stopped early at $2\times 10^{4}$ epochs when the testing error just begins to creep upward, these statistics tend even more strongly in its favor.
In this 1-to-1 comparison, we demonstrated that in a data generation and training regime borrowed from a well-cited Neural ODE work, our proposed KAN-ODE beats the performance of the Neural ODE in all key metrics. The apparent benefit of the Neural ODE’s faster iteration is rendered ineffective due to the KAN’s convergence to lower training and testing loss values in a short enough number of epochs to outscale the speedup of the MLP-based Neural ODE. We further study scaling effects in Fig. 3(C), where we present the converged training losses achieved by KAN-ODE and Neural ODEs of variable sizes. The KAN-ODEs larger than that studied above simply had wider hidden layers, while those smaller had narrower hidden layers as well as sparser grids. MLP-based Neural ODEs at each depth simply had more or fewer nodes in each hidden layer. Some larger Neural ODEs and smaller KAN-ODEs had their learning rates increased from the baseline value of $10^{-3}$ to avoid getting stuck in local minima. When comparing these KAN-ODEs and Neural ODEs of varying sizes, the benefit of the KAN-ODE becomes even more clear. Even with the modified RBF basis function, we appear to capture the $N^{-4}$ convergence rate of the third-order B-spline KAN proposed in [21], while the two MLP models appear to be converging at or even under an $N^{-2}$ rate. We see in fact that with even just the 240 KAN parameters studied in Figs. 3(A) and 3(B1) we are able to run the fourth-order convergence to saturation (larger KANs see no substantial performance gain), while traditional Neural ODEs would theoretically require upwards of thousands or even tens of thousands of parameters to reach similar performance, if they do not plateau earlier. The specific model sizes used in this convergence test are shown in Table 1, where the models used for Figs. 3(A) and 3(B1-B2) are in bold. Finally, we visualize the 80-parameter [2, 4, 4], [4, 2, 4] KAN-ODE’s activation functions in Fig. 3(D). We hypothesize based on certain similarities seen in these functions (i.e the three similar input activations coming from $x$), as well as the relatively sparse model of Eq. 11, that this KAN could be sparsified effectively for further interpretability. We further investigate sparse KANs and symbolic regression in Sec. III.2. In this example, we systematically explored the use of KAN-ODEs in a fairly simple ODE system and benchmarked their performance against comparable MLP-based Neural ODEs. We found that KAN-ODEs are an extremely efficient method to represent the predator-prey dynamics in this case, and beat Neural ODEs in all relevant metrics. In the next sections, we present larger-scale examples to showcase the flexibility and inference capabilities of KAN-ODEs.
III.2 Modeling Hidden Physics: Fisher-KPP PDE

### v2

@@ III Experiments
We present five inference examples in this work to demonstrate the broad capability of KAN-ODEs, grouped into three subsections. We first validated our methodology and compared its performance against the standard MLP-based Neural ODE in Sec. III.1, where the dynamics of the Lotka-Volterra predator-prey model are inferred. There we also comprehensively explored the neural scaling benefits of KAN-ODEs, as well as the efficient representations that can be inferred via sparsification and pruning. Then, we tackled a one-dimensional PDE in Sec. III.2, using a KAN-ODE to infer a symbolic source term in a wave propagation simulation. This example demonstrates the flexibility of KAN-ODEs to be used as submodels in higher complexity simulations, as well as their capability to directly extract symbolic models from data. Finally, we demonstrated the scalability and inference power of KAN-ODEs as standalone models in Sec. III.3 by inferring complete data-driven solutions to PDEs using limited samples. Examples here include shock wave formation governed by the Burgers’ equation and quantum wave function evolution governed by the Schrödinger equation, with an additional phase separation case using the Allen-Cahn equation provided in Appendix C.
III.1 KAN-ODEs vs Neural ODEs: Extensive Comparison via Lotka-Volterra
In this section, we targeted the Lotka-Volterra predator-prey model shown in Eq. 11. We replicated the simulation parameters used in the Neural ODE study of [9]. Namely, we used $\alpha=1.5$, $\beta=1$, $\gamma=1$, and $\delta=3$ with an initial condition of $\mathbf{u}_{0}=[x_{0},y_{0}]=[1,1]$ and a time span of $t\in[0,14]$ s. The first 3.5 s were used to train the KAN-ODE (Fig. 3(A)), while the remainder of the time history was withheld to validate the KAN-ODE’s ability to forecast the states at unseen times. The temporal grid used for training and loss evaluation had a spacing of 0.1 s. We tested the KAN-ODEs with various architectures shown in Table 1. The inputs to each KAN layer in this example were normalized to be on the $[-1,1]$ range required by the RBF networks via the hyperbolic tangent function, as in [32].
@@ III.1 KAN-ODEs vs Neural ODEs: Extensive Comparison via Lotka-Volterra
Loss profiles from training a 240-parameter KAN-ODE of shape [2, 10, 5], [10, 2, 5] as a representative case are shown in Fig. 3(B1). Strong convergence is observed at or even before $10^{4}$ epochs into training, while convergence into the $10^{-7}$ range of MSE loss occurs toward $10^{5}$ epochs. The converged KAN-ODE prediction is shown in Fig. 3(A), where the agreement in the testing window ($t\in[3.5,14]$) is nearly indistinguishable from the agreement in the training window ($t\in[0,3.5]$). We remark again that the trained KAN-ODE is reconstructing both the seen training data and unseen testing data given only the initial condition ($\mathbf{u}_{0}=[1,1]$), and can deliver output values on any temporal grid, thanks to its coupling to an ODE solver.
To contextualize these results, benchmark tests were performed to compare KAN-ODEs with standard, MLP-based Neural ODEs. We draw comparisons in terms of accuracy, convergence speeds, and neural scaling in Sec. III.1.1; and interpretability and generalization in Sec. III.1.2. In Sec. III.1.2, we additionally discuss the impacts of sparsified and symbolic KAN-ODEs.
III.1.1 Benchmark Tests and Neural Scaling Behavior
Figure 3: Comparison between KAN-ODE and Neural ODE for Lotka-Volterra predator-prey model. (A) Synthetic data (training and testing) and KAN-ODE reconstruction. (B) Loss profile during training for (B1) KAN-ODE and (B2) MLP-based Neural ODE of comparable sizes (240 and 252, respectively). (C) Comparison of converged KAN-ODE and Neural ODE error using different model sizes, and two MLP depths ($d=2$ and $d=3$). Neural scaling rates of $N^{-2}$ and $N^{-4}$ plotted for comparison, as per the theory in [26].
The MLP size and training procedure were taken from the Neural ODE work of [9] to avoid biasing the results in the KAN-ODE’s favor. Namely, an MLP with a hidden layer comprising 50 nodes and the hyperbolic tangent activation function replaced the KAN in Eq. 7. This MLP contained 252 trainable parameters (refer to the bold architecture in Table 1), which is comparable (and in fact slightly larger) than the 240 used in the reported KAN-ODE, facilitating fair comparison. Other relevant training details such as the train/test time windows, temporal grid, Lotka-Volterra equation parameters, and ODE solver remained identical to the KAN-ODE example, which itself was modeled as directly as possible after the Neural ODE efforts of [9].
Table 1: Lotka-Volterra tests with Neural ODEs and KAN-ODEs of varying size. Bold entries are studied further in Fig. 3.
@@ III.1.1 Benchmark Tests and Neural Scaling Behavior
In this 1-to-1 comparison using a data generation and training regime borrowed from a well-cited Neural ODE work, our proposed KAN-ODE beats the performance of the Neural ODE in all key metrics. Inspection of the loss profiles of Figs. 3(B1-B2) shows that the KAN-ODE converged to a substantially smaller training loss of $8.3\times 10^{-7}$ over $10^{5}$ epochs compared to the similarly-sized Neural ODE, which reached only $3\times 10^{-5}$ in the same number of epochs. The KAN-ODE qualitatively appears to overfit more than the Neural ODE model, especially in the later epochs. Quantitatively, however, it still beats the Neural ODE in accuracy to the unseen testing data. Even in the last 10% of the training cycle, i.e. epochs $90,000-100,000$, where the KAN-ODE appeared to overfit substantially, it still achieved an average testing loss of $6.8\times 10^{-3}$, with a minimum value in that range of $1.9\times 10^{-5}$. The Neural ODE in this same range averaged a much larger $1.4\times 10^{-2}$, with a minimum value of $5.4\times 10^{-5}$. If training for the KAN-ODE was stopped early at $2\times 10^{4}$ epochs when the testing error just begins to creep upward, these statistics would tend even more strongly in its favor.
In terms of computational cost, we found that the KAN-ODE iterated relatively slowly but actually converged 3-4x faster than the Neural ODE. On a single CPU core, the 240-parameter KAN-ODE took around 19 minutes per $10^{4}$ epochs, compared to the 252-parameter Neural ODE which took closer to 7 minutes per $10^{4}$ epochs. However, the KAN-ODE reached $2.6\times 10^{-5}$ training loss in just $10^{4}$ epochs, already beating the best-performing, $10^{5}$ epoch Neural ODE result of $3.0\times 10^{-5}$ training loss. Thus, while the KAN-ODE iterated 2.5 to 3x slower than the Neural ODE, it was able to reach the same predictive performance with 10x fewer epochs, resulting in an overall effective 3-4x speedup over the Neural ODE.
Figure 4: Sparsification, pruning, symbolic regression, and generalization results for Lotka-Volterra dynamics. (A) KAN with 72 parameters ([2, 3, 5], [3, 2, 5]), pruned from an initial 240 parameters ([2, 10, 5], [10, 2, 5]). (B) Symbolic KAN derived from (A), where each of the twelve activations is replaced by a univariate symbolic expression. (C) Generalization error when extrapolating outside of the yellow $(x,y)$ contour explored in the training and testing data. Each column represents a different gradient getter and has one row each for $dx/dt$ and $dy/dt$. Generalization is seen to improve when moving from MLP gradients to increasingly sparse KANs, and finally to symbolic representations fitted to KANs.
Up to this point, comparisons have been drawn exclusively between a single Neural ODE architecture and a single KAN-ODE architecture. Here, we discuss the effects of varied network architectures on the performance of KAN-ODEs and Neural ODEs, the results of which demonstrate a key scaling improvement of KAN-ODEs. Liu et al. [26] proposed that, thanks to the third-order splines used in the Kolmogorov-Arnold representations, KANs benefit from fourth-order neural scaling, or a quartic decrease in loss with respect to the KAN size (parameterized here as the number of trainable parameters). MLPs, on the other hand, often saturate quickly and can struggle to reach even first-order convergence [26, 40]. We investigated this phenomenon here by varying the layer width and grid size of KAN-ODEs, and by varying the layer width and depth of MLP-based Neural ODEs. Figure 3(C) shows that, even with the modified RBF basis function, we appear to capture the $N^{-4}$ B-spline KAN scaling rate with KAN-ODEs, while the two depths of MLP appear to be converging at at rate slower than $N^{-2}$. With just the 240 KAN parameters studied in Figs. 3(A) and 3(B1) we were able to run the fourth-order KAN-ODE convergence to saturation (larger KANs see no substantial additional performance gain), while MLP-based Neural ODEs would theoretically require upwards of thousands or even tens of thousands of parameters to reach similar performance, if they did not plateau earlier. The specific model sizes used in this convergence test are shown in Table 1, where the models used for Figs. 3(A) and 3(B1-B2) are in bold. To summarize, we found here that KAN-ODEs beat Neural ODEs in terms of data-fitting performance and computational efficiency in a systematic study of a broad range of network architectures and sizes.
III.1.2 Interpretation and Generalization of KAN-ODEs with Varying Sizes
In this section, we implemented sparsification, pruning, and symbolic regression to demonstrate a framework for developing interpretable KAN-ODE solvers. We additionally heuristically explored the generalizability of the various MLP and KAN-based models, where we found that KAN-ODEs have strong generalization potential even with limited training data. In addition to the 252-parameter Neural ODE and 240-parameter KAN-ODE as reference cases, three models were studied: a sparse KAN-ODE trained via parameter regularization and pruning, a symbolic KAN developed via symbolic regression on each univariate activation of the sparse KAN-ODE, and a symbolic representation for the entire model via symbolic regression on the global input-output pairs of the sparse KAN-ODE.
Sparsification, as proposed in [26] and discussed in [41], constitutes an additional term in the loss function to penalize non-zero trainable parameters, driving a given model toward sparser representations with fewer large activations. Here we use the simple L1 norm,
$\mathcal{L}\left(\bm{\theta}\right)=\text{MSE}\left(\mathbf{u}^{\text{KAN}}\left(t,\bm{\theta}\right),\mathbf{u}^{\text{obs}}\left(t\right)\right)+\gamma_{sp}|\bm{\theta}|_{1}.$ (12)
This is identical to Eq. 9, with the inclusion of the L1 penalty term with $\gamma_{sp}$ as a hyperparameter, which drives unnecessary activations toward zero. The logical next step is to reduce the size and complexity of the now-sparse KAN-ODE by eliminating any trivial nodes via pruning. To do so, we compare the training inputs and outputs of each node against a pruning parameter, $\gamma_{pr}$. Any node where the magnitude of at least one input or output exceeds $\gamma_{pr}$ is kept in the KAN-ODE, while those with all inputs and outputs smaller than $\gamma_{pr}$ are pruned. We began with a [2, 10, 5], [10, 2, 5] structure for a KAN (referred to henceforth as the dense KAN), based on observation from Fig. 3(C) of this architecture’s optimal performance. To sparsify the KAN-ODE, we included the regularization term of Eq. 12 with $\gamma_{sp}=5\cdot 10^{-4}$ in the training process. To prune, we trained this KAN-ODE to convergence, then eliminated nodes that did not satisfy $\gamma_{pr}=10^{-2}$ before training the remaining KAN-ODE to convergence once again. Figure 4(A) shows the sparse KAN-ODE with a final pruned size of [2, 3, 5], [3, 2, 5] trained using this procedure. Note that the darkness of the activation function curves is determined by the ratio of the outputs to the inputs (larger outputs have thicker lines).
Next, the symbolic KAN-ODE took the sparse KAN and fit symbolic expressions to each of its twelve univariate activation functions using the SymbolicRegression.jl package [42] with the four fundamental binary operators [+,,×,/][+,-$\times$,/][ + , - , × , / ]. Detailed results for each of these regressions is available in the Appendix A. Here, we plot the twelve symbolic activations for visualization in Fig. 4(B). To summarize, we first sparsified a 240-parameter, 10-node KAN into a 72-parameter, 3-node KAN. Then, we replaced these 72 parameters with just twelve univariate activation functions, a handful of which in the first layer are simple linear functions of the form $y=mx+b$. In terms of interpretability, we moved from a black box MLP-based Neural ODE to a visualizable yet bulky KAN, then to a visualizable and compact KAN, and finally to a straightforward collection of symbolic relationships that capture the Lotka-Volterra dynamics well. None of these steps required knowledge of the system outside of the provided training data.
Finally, we ran the entire Sparse KAN through SymbolicRegression.jl (rather than each activation separately) and found the following relationships which comprise the KAN representation,
$\frac{dx}{dt}$ $=1.495x-0.986xy,$ (13)
$\frac{dy}{dt}$ $=0.970xy-2.929y,$
which correspond nearly identically to our model parameters of 1.5, 1, 1, and 3. This result is effectively the pinnacle of interpretability and accuracy, though we believe is not a realistic option for most systems of larger-scale merit. Complete symbolic regression on an entire learned model or dataset is often not feasible, especially for complex models with higher dimensional inputs and outputs than are studied in this relatively straightforward ODE system [26]. Even when numerically manageable, clean convergence to compact functional forms is not guaranteed. On the other hand, a symbolic KAN like that shown in Fig. 4(B) requires symbolic regression only on a series of straightforward univariate activations and retains entirely interpretable intermediate steps. We present this global symbolic representation result as an interesting aside and a benchmark for comparison, but propose that sparse KAN-ODEs and symbolic KAN-ODEs are the most likely to have broader applicability in more realistic systems depending on the desired degree of interpretability and accuracy.
We move next to a discussion of the generalizability of KAN-ODEs. To conform to prior studies and faciliate fair comparison against MLP-based Neural ODEs, all models in this section were trained on data generated from a single initial condition, and tested on a continuation of these predator-prey oscillations for three more complete cycles (see Fig. 3). There exist many other potential $(x,y)$ combinations, however, and here we heuristically tested these trained models on their ability to generalize to completely unseen conditions.
The gradient error landscapes of the five models studied here (MLP-NODE, Dense KAN-ODE, Sparse KAN-ODE, Symbolic KAN-ODE, and symbolic representation of KAN-ODE) are shown from left to right, respectively, in Fig. 4(C), where the top row is the $dx/dt$ error and the bottom row is the $dy/dt$ error. The MLP-NODE results in the first column show near-zero error in the immediate vicinity of the training data contour, as expected from the performance metrics of Fig. 3. However, they also show significant gradient error in almost all ($x,y$) ranges not directly on or immediately near the training data contour, revealing the expected weak generalizability of the standard Neural ODE due to the fact that training data is limited to just those points on this contour.
The gradient error landscape of the dense KAN-ODE beats the MLP in generalizability, with much smaller gradient error values throughout the domain and especially inside of the yellow training contour where not only is the overall error substantially reduced, but we begin to additionally observe the formation of thin bands of near-perfect reconstruction away from the data contour. This result is rather interesting, as it was originally proposed in [26] that KANs beat MLPs in generalization thanks to their smaller number of parameters, but we observe here a substantial improvement with similarly sized models. In any case, these results suggest the potential of KAN-ODEs in the development of powerful models from limited data.
The gradient error landscape of the sparse KAN-ODE shows improvements over the dense KAN-ODE landscape. Inside the yellow data contour, we see further formation of near-perfect gradient reconstruction bands, surrounded by relatively low error regions. Results toward the edges of the landscape are mixed: the $dx/dt$ reconstruction captures the top left of the ($x,y$) domain with excellent accuracy but suffers from a large error in the upper right portion of the domain, while the $dy/dt$ reconstruction performs well toward low $y$ values, but suffers at high $y$ values. On the whole, we remark an overall improvement in generalization despite a nearly 4x reduction in KAN size, indicating the utility of sparse KAN-ODEs not just for compact and interpretable models, but also for improved generalization and extrapolation capabilities thanks to their smaller size. This result corroborates similar observations from [26] of the increase in generalizability with decreased KAN parameter counts.
The Symbolic KAN-ODE sees remarkable improvement in the $dx/dt$ generalization, with the low-error area being pushed out significantly toward high $x$ and $y$ values, further growth of the white bands of near-zero error throughout the domain, and a general lightening of the plot indicating better extrapolation and generalization. The $dy/dt$ reconstruction sees modest improvement, especially inside the data contour. We propose that this generalization improvement occurs due again to the more compact forms present in the symbolic KAN-ODE when compared to even the sparse KAN-ODE, and the extrapolation benefits that come with smaller system representations [26]. In the significantly out-of-training regions of $dx/dt$ where the symbolic KAN-ODE succeeds and the gridded basis functions of the dense KAN-ODE and Sparse KAN-ODE fail, we suggest that the input samples may be exiting the trained region of the RBF basis functions yet remaining within the useful range of the expressions used in the symbolic KAN-ODE.
The symbolic representation of the entire KAN obviously generalizes with near-perfect accuracy thanks to its discovery of what is essentially the complete underlying model. This result is plotted in the final column of Fig. 4(C) for reference, but we again emphasize that the other three KAN models are more reasonable choices for realistic applications: Dense KAN-ODEs succeeded with remarkably low training and testing error, sparse KAN-ODEs retained much of this low error while appearing to improve generalizability thanks to their smaller size, and Symbolic KAN-ODEs were substantially more interpretable and human-readable while also appearing to improve generalizability thanks to their smaller size and replacement of limited-scope basis functions with standard symbolic functions.
In this Lotka-Volterra example, we systematically and comprehensively explored the use of KAN-ODEs in a straightforward ODE system and benchmarked their performance against comparable MLP-based Neural ODEs in terms of accuracy, speed, model size, convergence rate, interpretability, and generalization. We found that KAN-ODEs are an extremely efficient method to represent the predator-prey dynamics in this case, and beat Neural ODEs in all relevant quantitative metrics. We additionally explored the interpretability of KAN-ODEs via symbolic regression of each activation function, and then heuristically evaluated a series of MLP and KAN-based ODE approaches in terms of generalizability to unknown data, where we again found KAN-ODEs superior. In the next sections, we present larger-scale examples to showcase the flexibility and inference capabilities of KAN-ODEs.
III.2 Modeling Hidden Physics in PDEs: Fisher-KPP PDE